# A pedagogical approach to the black hole information issue

### v3

@@ Abstract
We provide a pedagogical introduction to the concepts underlying black hole information loss, intended for readers familiar with special relativity and quantum mechanics. We emphasize that there is no paradox of information loss, and that proposals suggesting deviations from well-established theories at arbitrary regimes are inherently contradictory.
@@ I Introduction
The concept of a region from which nothing can escape due to gravity, known as a black hole, has been a notable front for developments in theoretical physics over the last decades. One of the most distinguished results from these developments is the state of affairs commonly referred to as the black hole information paradox. This unfortunate nomenclature follows mainly from an erroneous interpretation of the implications black holes have for information preservation, which has led to many physically unreliable proposals. In order to fully grasp the nature of this issue and understand associated the misconceptions, it is useful to begin with a review of the current descriptions of the fundamental interactions.
General relativity (GR) Wald1984 is a theory of the gravitational interaction, whose characterization is made by the curvature of spacetime due to the presence of an energy distribution. A qualitative notion of this description is given in fig. 1, inspired by the quote (Misner1973, , p. 5): “Spacetime tells matter how to move; matter tells spacetime how to curve”.
@@
This classical [^1] By classical, it is meant that the characterization of the gravitational interaction, given by the spacetime geometry, assumes definite values in any length scale, not only those much larger than atomic ones (i.e., distances of order $10^{-9}\;\text{m}$). theory currently paints the most accurate picture of gravitational phenomena, and thus, is the one we use to describe black holes. On the other hand, the present depiction of the electroweak and strong interactions is given by quantum field theory (QFT) Parker2009 , which can be understood as an extension of quantum mechanics (QM) and special relativity (SR) that treats quantized fields as protagonists of these interactions. Being highly successful in describing observational data and experiments in both of their pertinent regimes of application, these theories can be combined in the mathematically rigorous, but physically incomplete, formalism of quantum field theory in curved spacetime (QFTCS).
The argument behind the physical incompleteness of QFTCS follows from GR, which quantifies how a classical energy distribution, described by the energy-momentum tensor, affects the spacetime geometry. In QFTCS, the quantum field dynamics is governed by the curved spacetime, but one does not know how to generally describe the effects of the quantum field energy content on the spacetime geometry, known as backreaction effects (see fig. 2).
Figure 2: Qualitative description of the interactions in QFTCS. The dashed line represents the fact that in this framework there is no general description of backreaction effects.
Nonetheless, one can rely on physical arguments to qualitatively describe possible backreaction for specific phenomena, which leads to the more complete framework of semiclassical gravity (SG) [^2] Although we distinguish between QFTCS and SG due to the consideration of backreaction effects, some references use these terms as synonyms.. Consequently, following the experimental success of GR and QFT, SG provides the best available description of the fundamental interactions. In other words, SG provides the best formalism available to study quantum effects in and on gravitation.
A notable quantum effect is the Hawking effect, which states that stationary (i.e., time independent) black holes should effectively emit [^3] Hawking radiation does not “get out” from a black hole. an approximate thermal spectrum of particles, known as Hawking radiation. Now, if one considers the implication of a negative energy flux entering the black hole due to local energy-momentum conservation, then the main backreaction effect due to Hawking radiation will be to reduce the black hole mass and other physical parameters, resulting in an effective evaporation process. Hence, after a finite time, a black hole will reach the Plank scale, in the sense that its radius will be of order of the Planck length,
@@
To understand how black holes lead to information loss in SG, consider the following line of reasoning. Suppose one starts with a stationary energy distribution, which is described by its mass, angular momentum, electric charge, baryon and lepton number, and any other physical property. From the perspective of QM, complete knowledge of the quantum state of this distribution would correspond to a pure state. An example of this would be a highly diluted cold matter distribution (e.g., a dust cloud) such that one has all the possible information about its configuration. Of course, such a distribution would not be exactly stationary due gravitation and other interactions, but, in distant regions it would be a good approximation to describe it by a stationary spacetime. Now, as this distribution evolves over time, it could form, say, a star. If such a star is massive enough, at the later stages of its life gravitational effects will not be sustained, and a black hole will form as a consequence of gravitational collapse. Even though the black hole formation process is expected to be highly dependent on the energy distribution details and how it collapses, asymmetries of the distribution are expected to be radiated away in the form of gravitational waves. Following a conjecture that black holes must eventually reach a stationary configuration, the initial energy distribution will then give rise to a black hole described only by three parameters: its mass, angular momentum, and electric charge. However, the Hawking effect predicts that such black holes should behave as a gray body, effectively emitting an approximately thermal spectrum of radiation. This radiation carries away the mass, angular momentum, and electric charge of the black hole, and, after a finite time, the black hole might completely evaporate, leaving only radiation. In light of SG, the resulting quantum state would necessarily correspond to a mixed one, which arises in a context in which one does not have all the possible information about the state. Therefore, the complete evaporation of a black hole would take a pure state to a mixed one, corresponding to the fact that most information about the details of the energy distribution that gave rise to the black hole is lost.
@@
Evidently, in order for this conclusion to hold, one would have to trust SG predictions at arbitrary scales. Indeed, even though SG predicts an evaporation process for regimes far from the Planck scale, when a black hole reaches a radius of order $\ell_{p}$, only a theory of quantum gravity would be appropriate to accurately describe its fate. Consequently, it is an interesting open question how a theory of quantum gravity would address black hole evaporation. Nonetheless, the conclusion of information loss in the framework of SG contradicts no known phenomena and established theories; in fact, it is a physical prediction. In contrast, it has been extensively and incorrectly argued in the literature that the conclusion of information loss is paradoxical in light of unitary evolutions in QM. As a result, many inconsistent proposals have come to light to solve a nonexistent paradox.
@@
The purpose of this paper is twofold. First, we provide a pedagogical presentation of the concepts that lead to the conclusion of information loss due to black holes in SG. Second, we show that there is no paradox of information loss, and discuss the flaws in proposals that seek to invalidate SG predictions at arbitrary regimes. The paper is organized as follows. In Sec. II, we discuss black holes in the framework of GR, introducing the necessary ideas for their definition and analysis. In Sec. III, we present the Hawking effect and the implication of quantum entanglement for black holes in SG. In Sec. IV, we show how SG predicts that black holes will result in information loss, and discuss the misconceptions surrounding this conclusion. We conclude in Sec. V with the final remarks. We assume familiarity with SR and QM. The arguments presented are discussed in much more detail in Bergamaschi2024 , which includes an extensive pedagogical review of the technical analysis necessary to support the results stated in this paper. We work with SI units, but we set $c=1$ for all spacetime diagrams simply to allow for more intuitive illustrations. Unless stated otherwise, we will consider four-dimensional spacetimes.
@@ II Classical aspects of black holes
In the purely classical framework of GR, a black hole is defined as a region of spacetime from which nothing can escape. In this section, we will detail the classical concepts necessary to make this notion precise and explain the importance of stationary black holes. Our starting point is to discuss pertinent properties of Minkowski spacetime.
@@ II.1 Minkowski spacetime
In SR, the background for all phenomena is the Minkowski spacetime, which is a collection of events endowed with a four-dimensional measure (i.e., the invariant interval) to every pair of such events given by the Minkowski metric. The concept of causality follows from the possible events that can influence one another, either by a massive particle or an electromagnetic wave. The analysis of which events can be causally connected to one another then defines the causal structure of Minkowski spacetime, which is given by the set of events that can be interpreted as the future and past of events.
For instance, consider an event in Minkowski spacetime, $a$. Associated to $a$ are two sets of events: those that can be connected to it by a timelike or null curve [^4] These curves must obey pertinent differentiability conditions and be oriented in a single direction of time (see (Bergamaschi2024, , sec. 2.2) for more details)., and those that can be connected to it only by a spacelike curve. The latter is the set of events that are causally disconnected from $a$, while the former is the set of events that are causally connected to $a$. In this manner, the notion of future and past for causally connected events arises from the fact that either a massive particle or an electromagnetic wave can be emitted to influence $a$ or to have emanated from $a$. These ideas are illustrated in fig. 3. Notice that any event inside the gray region is causally connected to $a$, as well as those along the null curves that pass through $a$. Events not in the gray regions or on its boundary are causally disconnected from $a$, meaning that only a spacelike curve can connect them to $a$ (i.e., faster than light travel).
@@
In particular, the distinction between the future and past sets is given by a predetermined continuous definition of what constitutes going “forwards” or “backwards” in time. In this manner, all curves considered in the following should be understood as “going” from the past to the future.
A helpful way to visualize the entire causal structure of Minkowski spacetime is by a conformal diagram (also known as a Carter-Penrose diagram). Such a diagram allows one to depict the entire spacetime in a finite illustration, while keeping null curves aligned at $\pm 45\degree$. The conformal diagram of Minkowski spacetime is illustrated in fig. 4. In the coordinate system $\{(t,r,\theta,\phi)\}$, where $t$ is the time parameter of static inertial observers in the usual spherical coordinates $\{(r,\theta,\phi)\}$ , the interpretation of the conformal diagram is as follows [^5] In the following, all conformal diagrams will be presented with two angular coordinates suppressed.. The vertical black line is the center of the coordinate system, $r=0$, for different values of $t$. All null curves begin at the three-dimensional null hyperplane $\mathscr{I}^{-}$ and end at three-dimensional null hyperplane $\mathscr{I}^{+}$. Similarly, all timelike curves begin at the point $\mathcal{I}^{-}$ and end at the point $\mathcal{I}^{+}$. Additionally, the red line is a three-dimensional spacelike section (e.g., a $t=\text{constant}$ hyperplane) that begins at $r=0$ and ends at $\mathcal{I}^{0}$. A diagram analogous to that of fig. 3 is also represented in fig. 4, where the null curves emanating from a two-sphere, $\mathscr{A}$, should be understood to extend until they reach $\mathscr{I}^{-}$ and $\mathscr{I}^{+}$. In this manner, the future and past of any other two-sphere can be readily verified by tracing null curves through it, i.e., delimiting its light cone.
@@
One of the most notable features of the conformal diagram of Minkowski spacetime are the two null hyperplanes $\mathscr{I}^{+}$ and $\mathscr{I}^{-}$. These are known as future null infinity and past null infinity, respectively. They can be interpreted as “infinitely distant regions” of Minkowski spacetime such that only light rays can reach, to the future and to the past. This is due to the fact that all null curves begin at $\mathscr{I}^{-}$ and end at $\mathscr{I}^{+}$. To visualize this, consider a light ray that was emitted far away from the origin of the coordinate system (i.e., $r\gg 0$) at early times (i.e., $t\ll 0$). If the light ray was emitted so that its radial coordinate decreases over time (i.e., dr/dt=cdr/dt=-c), then it will eventually reach the origin of the coordinate system, $r=0$. After that, its radial coordinate will increase over time (i.e., $dr/dt=c$), and it will eventually reach the regions far away from the origin at late times (i.e., $t\gg 0$). The regions $\mathscr{I}^{-}$ and $\mathscr{I}^{+}$ appear in this line of reasoning when one takes the limit of $r\to\infty$ (i.e., far away region) and $t\to\pm\infty$ (i.e., early and late times). In particular, the process of the light ray reaching $r=0$ and thus being described by a different null curve is illustrated by the “reflection” of the null curve in fig. 4 (see also fig. 5 for more examples).
To conclude our analysis of Minkowski spacetime, we mention an important property of its causal structure. From its conformal diagram, it can be deduced that any three-dimensional spacelike section that begins at $r=0$ and extends to $\mathcal{I}^{0}$ will be intersected exactly once by all null curves. More specifically, consider two three-dimensional spacelike sections that begin at $r=0$ and extend to $\mathcal{I}^{0}$, denoted by $\Sigma_{0}$ and $\Sigma_{1}$. From fig. 5, it is possible to see that all null curves emerge from $\mathscr{I}^{-}$ and intersect $\Sigma_{0}$ and $\Sigma_{1}$ exactly once, either on their way to reach $r=0$ or on their way to reach $\mathscr{I}^{+}$ after “reflecting” at $r=0$. Because of this property, $\Sigma_{0}$ and $\Sigma_{1}$ are referred to as Cauchy surfaces. Spacetimes possessing a Cauchy surface are said to be globally hyperbolic, which can be interpreted as those that obey pertinent conditions on causality and determinism, allowing one to formulate a well-posed problem for dynamics.
@@
Additionally, globally hyperbolic spacetimes can be described in terms of a foliation of Cauchy surfaces, i.e., they can be written as a collection of Cauchy surfaces. An example of this is a foliation of Minkowski spacetime in terms of $t=\text{constant}$ hyperplanes. Although a foliation is not unique, in a given foliation each Cauchy surface can be interpreted as an “instant of time” of the spacetime. In this sense, information on a Cauchy surface can be used to determine the entire development of the spacetime, i.e., physical conditions on any other Cauchy surface. Indeed, because each null curve must intercept a Cauchy surface exactly once, information can be interpreted to be carried over (i.e., conserved) from one Cauchy surface to the others. This can be exemplified by considering both $\Sigma_{0}$ and $\Sigma_{1}$, as all null curves that intersect $\Sigma_{0}$ will eventually intersect $\Sigma_{1}$. Conversely, complete predictability from physical conditions at some “instant of time” will not occur for spacetimes that fail to be globally hyperbolic. Consequently, it is challenging to envision how information could be conserved when one is analyzing non-globally hyperbolic spacetimes.

### v4

@@ Abstract
We provide a pedagogical introduction to the concepts underlying black hole information loss, intended for readers familiar with special relativity and quantum mechanics. We emphasize that there is no paradox of information loss and that proposals that deviate from well-established theories in arbitrary regimes are inherently contradictory.
@@ I Introduction
The concept of a region from which nothing can escape due to gravity, known as a black hole, has been a notable front for developments in theoretical physics over the last few decades. One of the most remarkable outcomes of these developments is the state of affairs commonly referred to as the black hole information paradox. This unfortunate nomenclature stems mainly from an erroneous interpretation of the implications of black holes for information preservation, which has led to many physically unreliable proposals. To fully grasp the nature of this issue and understand the associated misconceptions, it is beneficial to begin with a review of the current descriptions of the fundamental interactions.
General relativity (GR) Wald1984 is a theory of the gravitational interaction whose characterization is made by the curvature of spacetime due to the presence of an energy distribution. A qualitative notion of this description is given in fig. 1, inspired by the quote (Misner1973, , p. 5): “Spacetime tells matter how to move; matter tells spacetime how to curve”. This classical [^1] By classical, it is meant that the characterization of the gravitational interaction, given by the spacetime geometry, assumes definite values in any length scale, not only those much larger than atomic ones (i.e., distances of order $10^{-9}\;\text{m}$). theory currently paints the most accurate picture of gravitational phenomena, and thus, is the one we use to describe black holes. On the other hand, the present depiction of the electroweak and strong interactions is given by quantum field theory (QFT) Parker2009 , which can be understood as an extension of quantum mechanics (QM) and special relativity (SR) that treats quantized fields as protagonists of these interactions. Being highly successful at describing observational data and experiments in both their pertinent regimes of application, these theories can be combined in the mathematically rigorous but physically incomplete formalism of quantum field theory in curved spacetime (QFTCS).
@@
The argument behind the physical incompleteness of QFTCS follows from GR, which quantifies how a classical energy distribution, described by the energy-momentum tensor, affects the spacetime geometry. In QFTCS, the quantum field dynamics is governed by the curved spacetime, but one does not know how to generally describe the effects of the quantum field energy content on the spacetime geometry, known as backreaction effects (see fig. 2). Nonetheless, one can rely on physical arguments to qualitatively describe possible backreaction for specific phenomena, which leads to the more complete framework of semiclassical gravity (SG) [^2] Although we distinguish between QFTCS and SG due to the consideration of backreaction effects, some references use these terms as synonyms.. Consequently, following the experimental success of GR and QFT, SG provides the best available simultaneous description of the fundamental interactions. In other words, SG provides the best formalism available to study quantum effects in and on gravitation.
Figure 2: Qualitative description of the interactions in QFTCS. The dashed line indicates that, within this framework, there is no general description of backreaction effects.
A notable quantum effect is the Hawking effect, which states that stationary (i.e., time independent) black holes effectively emit [^3] Hawking radiation does not “get out” from a black hole. an approximate thermal spectrum of particles, known as Hawking radiation. Now, if one considers the implication of a negative energy flux entering the black hole due to local energy-momentum conservation, then the main backreaction effect due to Hawking radiation will be to reduce the black hole mass and other physical parameters, resulting in an effective evaporation process. Hence, after a finite time, a black hole will reach the Planck scale, in the sense that its radius will be of the order of the Planck length,
@@
To understand how black holes lead to information loss in SG, consider the following line of reasoning. Suppose one starts with a stationary energy distribution, which is described by its mass, angular momentum, electric charge, baryon and lepton number, and any other physical property. From the perspective of QM, complete knowledge of the quantum state of this distribution would correspond to a pure state. An example of this would be a highly diluted cold-matter distribution (e.g., a dust cloud) such that one has all the possible information about its configuration. Of course, such a distribution would not be exactly stationary due to gravitation and other interactions, but, in distant regions, it would be a good approximation to describe it by a stationary spacetime. Now, as this distribution evolves, it could form, say, a star. If such a star is sufficiently massive, at the later stages of its life, gravitational effects will not be sustained, and a black hole will form as a consequence of gravitational collapse. Even though the black hole formation process is expected to be highly dependent on the energy distribution details and how it collapses, asymmetries of the distribution are expected to be radiated away in the form of gravitational waves. Following a conjecture that black holes must eventually reach a stationary configuration, the initial energy distribution will then give rise to a black hole described only by three parameters: its mass, angular momentum, and electric charge. However, the Hawking effect predicts that such black holes should behave as gray bodies, effectively emitting an approximate thermal spectrum of radiation. This radiation carries away the black hole mass, angular momentum, and electric charge, and after a finite time, the black hole might completely evaporate, leaving only radiation. In light of SG, the resulting quantum state would necessarily correspond to a mixed one, which arises in a context in which one does not have all the possible information about the state. Therefore, the complete evaporation of a black hole would take a pure state to a mixed one, corresponding to the fact that most information about the details of the energy distribution that gave rise to the black hole is lost.
@@
Evidently, for this conclusion to hold, one would have to trust SG predictions at arbitrary scales. Indeed, even though SG predicts an evaporation process for regimes far from the Planck scale, when a black hole reaches a radius of order $\ell_{p}$, only a theory of quantum gravity would be appropriate to describe its fate accurately. Consequently, it is an interesting open question how a theory of quantum gravity would address black hole evaporation. Nonetheless, the conclusion of information loss in the framework of SG contradicts no known phenomena and established theories; in fact, it is a physical prediction. In contrast, it has been extensively and incorrectly argued in the literature that the conclusion of information loss is paradoxical in light of unitary evolutions in QM. As a result, many inconsistent proposals have come to light to solve a nonexistent paradox.
@@
The purpose of this paper is twofold. First, we provide a pedagogical presentation of the concepts that lead to the conclusion that information is lost due to black holes in SG. Second, we show why there is no paradox of information loss and discuss the flaws in proposals that seek to invalidate SG predictions at arbitrary regimes. The paper is organized as follows. In sec. II, we discuss black holes in the framework of GR, introducing the necessary ideas for their definition and analysis. In sec. III, we present the Hawking effect and the implication of quantum entanglement for black holes in SG. In sec. IV, we show how SG predicts that black holes lead to information loss and discuss misconceptions surrounding this conclusion. We conclude in sec. V with the final remarks. We assume familiarity with SR and QM. The arguments presented are discussed in much more detail in Bergamaschi2024 , which includes an extensive pedagogical review of the technical analysis required to support the ideas of this paper. We work with SI units, but we set $c=1$ for all spacetime diagrams [^4] This is a mere visual simplification to allow null curves to be aligned with lines of slope $\pm 1$.. Unless stated otherwise, we will consider four-dimensional spacetimes.
@@ II Classical aspects of black holes
In the purely classical framework of GR, a black hole is defined as a region of spacetime from which nothing can escape. In this section, we detail the classical concepts necessary to make this notion precise and explain the importance of stationary black holes. Our starting point is to discuss pertinent properties of Minkowski spacetime.
@@ II.1 Minkowski spacetime
In SR, the background for all phenomena is the Minkowski spacetime, which is a collection of events endowed with a four-dimensional measure (i.e., the invariant interval) to every pair of such events given by the Minkowski metric. The concept of causality follows from the possible events that can influence one another, say, by a massive particle or an electromagnetic wave. The analysis of which events can be causally connected then defines the causal structure of Minkowski spacetime, which is given by the set of events that can be interpreted as the future and past of events.
For instance, consider an event in Minkowski spacetime, $a$. Associated to $a$ are two sets of events: those that can be connected to it by a timelike or null curve [^5] These curves must obey pertinent differentiability conditions and be oriented in a single direction of time (see (Bergamaschi2024, , sec. 2.2) for more details)., and those that cannot. The latter is the set of events that are causally disconnected from $a$, while the former is the set of events that are causally connected to $a$. In this manner, the notion of future and past for causally connected events arises from the fact that either a particle or wave can be emitted to influence $a$ or to have emanated from $a$. These ideas are illustrated in fig. 3. Note that any event inside the gray region is causally connected to $a$, as well as those along the null curves that pass through $a$. Events not in the gray regions or on its boundary are causally disconnected from $a$, meaning that only faster-than-light travel would allow one to reach them from $a$. In particular, the distinction between the future and past sets is given by a predetermined continuous definition of what constitutes going “forwards” or “backwards” in time. All curves considered in the following should be understood as “going” from the past to the future.
@@
A helpful way to visualize the entire causal structure of Minkowski spacetime is by a conformal diagram (also known as a Carter-Penrose diagram). Such a diagram allows one to depict the entire spacetime in a finite illustration, while keeping null curves aligned with lines of slope $\pm 1$. The conformal diagram of Minkowski spacetime is illustrated in fig. 4. In the coordinate system $\{(t,r,\theta,\phi)\}$, where $t$ is the time parameter of static inertial observers in the usual spherical coordinates $\{(r,\theta,\phi)\}$, the interpretation of the conformal diagram is as follows [^6] In the following, all conformal diagrams will be presented with two angular coordinates suppressed.. The vertical black line is the center of the coordinate system, $r=0$, for different values of $t$. All null curves begin at the three-dimensional null hyperplane $\mathscr{I}^{-}$ and end at the three-dimensional null hyperplane $\mathscr{I}^{+}$. Similarly, all timelike curves begin at the point $\mathcal{I}^{-}$ and end at the point $\mathcal{I}^{+}$. Additionally, the red line is a three-dimensional spacelike section (e.g., a $t=\text{constant}$ hyperplane) that begins at $r=0$ and ends at $\mathcal{I}^{0}$. A diagram analogous to that of fig. 3 is also represented in fig. 4, where the null curves emanating from the two-sphere, $\mathscr{A}$, should be understood to extend until they reach $\mathscr{I}^{-}$ and $\mathscr{I}^{+}$. Consequently, the future and past of any other two-sphere can be readily verified by tracing null curves through it, i.e., delimiting its light cone.
@@
One of the most notable features of the conformal diagram of Minkowski spacetime are the two null hyperplanes $\mathscr{I}^{+}$ and $\mathscr{I}^{-}$. These are known as future null infinity and past null infinity, respectively. They can be interpreted as “infinitely distant regions” of Minkowski spacetime, such that only light rays can reach, to the future and to the past (i.e., asymptotic future and past). This is because all null curves begin at $\mathscr{I}^{-}$ and end at $\mathscr{I}^{+}$. To visualize this, consider a light ray that was emitted far away from the origin of the coordinate system (i.e., $r\gg 0$) at early times (i.e., $t\ll 0$). If the light ray was emitted so that its radial coordinate decreases over time (i.e., dr/dt=cdr/dt=-c), then it will eventually reach the origin of the coordinate system, $r=0$. After that, its radial coordinate will increase over time (i.e., $dr/dt=c$), and it will eventually reach the regions far away from the origin at late times (i.e., $t\gg 0$). The regions $\mathscr{I}^{-}$ and $\mathscr{I}^{+}$ appear in this line of reasoning when one takes the limit of $r\to\infty$ (i.e., far away region) and $t\to\pm\infty$ (i.e., early and late times). In particular, the process of the light ray reaching $r=0$ and thus being described by a different null curve is illustrated by the “reflection” of the null curve in fig. 4 (see also fig. 5 for more examples).
To conclude our analysis, we mention an important property of Minkowski spacetime: it possesses a Cauchy surface [^7] A Cauchy surface is a closed hypersurface which is intersected by each inextendible timelike and null curve exactly once (see (Bergamaschi2024, , sec. 2.2) for more details).. To understand this in more detail, consider any three-dimensional spacelike section that begins at $r=0$ and extends to $\mathcal{I}^{0}$, such as $\Sigma_{0}$ and $\Sigma_{1}$ in fig. 5. It is possible to see that all timelike and null curves intersect $\Sigma_{0}$ and $\Sigma_{1}$ exactly once. For instance, null curves emerge from $\mathscr{I}^{-}$ and intersect $\Sigma_{0}$ and $\Sigma_{1}$ exactly once, either on their way to reach $r=0$ or on their way to reach $\mathscr{I}^{+}$ after “reflecting” at $r=0$. Thus, both $\Sigma_{0}$ and $\Sigma_{1}$ are Cauchy surfaces. Spacetimes possessing a Cauchy surface are said to be globally hyperbolic, which can be interpreted as those that obey pertinent conditions on causality and determinism, allowing one to formulate a well-posed problem for dynamics.
@@
Globally hyperbolic spacetimes can be described as a foliation of Cauchy surfaces, i.e., as a collection of Cauchy surfaces. An example of this is a foliation of Minkowski spacetime by $t=\text{constant}$ hyperplanes. Although a foliation is not unique, in a given foliation, each Cauchy surface can be interpreted as an “instant of time” of spacetime. Essentially, information on a Cauchy surface can be used to determine the entire development of spacetime, i.e., physical conditions on any other Cauchy surface. Indeed, because each null curve must intercept a Cauchy surface exactly once, information can be interpreted to be carried over (i.e., conserved) from one Cauchy surface to the others. This can be exemplified by considering both $\Sigma_{0}$ and $\Sigma_{1}$, as all null curves that intersect $\Sigma_{0}$ will eventually intersect $\Sigma_{1}$. Conversely, complete predictability from physical conditions at some “instant of time” will not hold for spacetimes that are not globally hyperbolic. Consequently, it is challenging to envision how information could be preserved when analyzing non-globally hyperbolic spacetimes.